Kyle is awake exactly 18 hours/day. While he is awake, Kyle either works at The Adjustment Bureau, which pays him $12.50/hour, or he just hangs out which pays nothing. Kyle spends whatever money he earns at We-Can-Remember-It-For-You-Wholesale Travel Agency which charges $15 to implant ("one unit of") false, but totally convincing, memories of great vacations in Kyle's brain. Draw Kyle's "budget line" showing all the combinations of hang-out-time and the amount of false memory he can afford. Be sure to measure the hang-out-time on the horizontal axis. What is the opportunity cost of hanging out one extra hour measured in terms of the amount of false memory he would have to give up?
Answer #1

Let the number of hang out hours=H

Let the number of units of false memory purchased=F

Wage rate=w=$12.50

Price of one unit of false memory=r=$15

Labor hours=(18-H)

Labor income=(18-H)*12.50=225-12.50H

Labor income is used in purchasing false memory. So,

Labor income=total purchase amount of false memory

225-12.5H=15*F

F+(5/6)*H=15 (budget line)

On rearranging we get

F=-(5/6)*H+15  

Opportunity Cost of hanging out an extra hour=slope of above line=5/6 units of false memory

H F=-(5/6)*H+15
0 15
3 12.5
6 10
9 7.5
12 5
15 2.5